INTERSTELLAR H_{3}^{+} IN METASTABLE ROTATIONAL LEVELS

\underline{TAKESHI OKA}, Department of Chemistry and Department of Astronomy and Astrophysics, the Enrico Fermi Institute, the University of Chicago, Chicago, IL 60637.

A recent tentative detection of interstellar H_{3}^{+} toward the Galactic center has revealed the possible presence of H_{3}^{+} in a metastable rotational level (J, K) = (3, 3).^{a} The absence of H_{3}^{+} in the lower (2, 2) and (2, 1) rotational levels during the same observation demonstrates a non-thermal populational distribution in which H_{3}^{+} in the two lower levels relax to the lowest (1, 1) level by spontaneous emission due to a dipole moment induced by centrifugal distortion,^{b} while those in the metastable (3, 3) level do not relax radiatively.^{c} Such metastability is expected to occur also for rotational levels with J = K = 5, 6, 7 etc. and a few other levels. H_{3}^{+} in the metastable levels thermalize only through reactive collisions with other species, mainly H_{2}, H, He and electrons. We discuss thermalization of H_{3}^{+} and the resulting non-thermal population in the interstellar medium taking into account both the spontaneous emission and the collision and the production and destruction of H_{3}^{+}. Properly interpreted, H_{3}^{+} in metastable rotational levels will provide us with a useful astrophysical probe to measure temperature and density of the interstellar medium with relatively high temperature.

aM. Goto, B. J. McCall, T. R. Geballe, T. Usuda, N. Kobayashi, H. Terada and T. Oka, PASJ \underline{54}, 951 (2002)

bJ. K. G. Watson, J. Mol. Spectrosc. \underline{40}, 536 (1971)

cF. S. Pan and T. Oka, ApJ \underline{305}, 518 (1986)